OpenStudy (anonymous):

What is the volume of this right circular cylinder? The radius is 4 cm and the height is 6 cm.

OpenStudy (anonymous):

pi*r^2*h = pi*(4)^2*(6) = 301.5928 cm^3

OpenStudy (anonymous):

Thanks :)